USING A NON-PRODUCT QUADRATURE GRID TO COMPUTE THE VIBRATIONAL SPECTRUM OF C_{2}H_{4}

\underline{GUSTAVO AVILA} and TUCKER CARRINGTON JR., Chemistry Department, Queen’s University, Kingston, Ontario K7L 3N6, Canada.

We present an accurate 12-D basis set calculation of the lowest 100 energy levels of the C_{2}H_{4} molecule. A Smolyak non-product quadrature grid, a pruned product basis set, and the Lanczos algorithm are used. This scheme allows one to reduce the size of the basis set by almost 7 orders of magnitude (from 9 × 10^{12} to 1.3 × 10^{6}) and the size of the quadrature grid by almost 6 orders of magnitude (from 5.6×10^{13} to 1.52 × 10^{8}). Basis pruning and the nonproduct quadrature grid therefore enable us to solve a problem, numerically exactly, that would be impossible without these tools.